OpenStudy (anonymous):

Find the solution to the equation log16 32=x+5. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

rewrite usine exponential: \[\log_{16}32=x+5 \rightarrow 16^{x+5}=32 \rightarrow \left( 2^{4} \right)^{x+5}=2^{5}\rightarrow 2^{4\left( x+5 \right)}=2^{5}\]

OpenStudy (anonymous):

so the total is 32? @pgpilot326

OpenStudy (anonymous):

how would you explain the procedure?

OpenStudy (anonymous):

you are rewriting the log as an exponential equation, then changing each term to a common base and then equating the exponents and solving the resulting equation for the unknown.

OpenStudy (anonymous):

Thank you, I appreciate the help!

OpenStudy (anonymous):

do you understand?